Chapter 11 (11.45)

Ethanol (C2H5OH) melts at -114 ∘C and boils at 78 ∘C. The enthalpy of fusion of ethanol is 5.02 kJ/mol, and its enthalpy of vaporization is 38.56 kJ/mol. The specific heat of solid and liquid ethanol are 0.97 J/g⋅K are 2.3 J/g⋅K respectively.

How much heat is required to convert 22.0 g of ethanol at -152 ∘C to the vapor phase at 78 ∘C?
1
Expert's answer
2016-11-09T08:57:10-0500
Q=0,97*(-114-(-152))*22+2.3*(78-(-114))*22+5020*22/(2*12+6+16)+38560*22/(12*2+6+16)=810,92+9715,2+2400,87+18441,74=31368,73J